# UniBEV: Multi-modal 3D Object Detection with Uniform BEV Encoders for Robustness against Missing Sensor Modalities

### v2

@@ IV-B Multi-Modal 3D Object Detection
Table I showcases the inference performance of the fusion-based detector on both multi-modal input, as well as single-modality input using the same trained weights. The summary metrics of our UniBEV ($58.7\%$ NDS and $52.5\%$ mAP) significantly surpass the baseline methods, indicating UniBEV is more robust over varying input modalities.
On all input modalities UniBEV performs better than its multi-modal baselines, achieving $68.5\%$ NDS and $64.2\%$ mAP for LiDAR+camera fusion, especially the difference in camera-only performance is notable. Despite employing a more powerful image backbone [36], BEVFusion lags markedly behind UniBEV when only cameras input is available.
Given that the CenterPoint head of BEVFusion and our detection head exhibit comparable detection capabilities (as evidenced by the near identical performance of UniBEV_L and CenterPoint), the performance difference for camera-only between UniBEV and BEVFusion can be attributed to the quality of BEV features and its fusion strategy. Fig. 3 illustrates that compared to BEVFusion, UniBEV’s camera and LiDAR BEV features more clearly discern similar object locations, and that these are better spatially aligned. Besides, BEVFusion utilizes LSS [6] as the camera BEV encoder to project image features into the BEV feature space. This enforces an inductive bias on its camera BEV features not present in its LiDAR BEV features, as exhibited by the hexagon-shaped outline.
Figure 3: Example of the BEV feature maps of different modalities for a single sample. High intensity indicates high variance across channels at that location. UniBEV aligns modalities for object detection, resulting in strong responses at the same locations in both camera and LiDAR.
@@
To support the comparison between UniBEV and its multi-modal baselines, we show some qualitative detection results in Fig. 4. For instance, we can see that BEVFusion for camera-only suffers from various false negatives, whereas MetaBEV tends to have more false positives.
@@
Fig. 5 (a) visualizes UniBEV’s $N=256$ normalized CNW weights, sorted from most LiDAR weighted to most camera weighted, and also reports their total summed weights. We observe the sum of camera weights is smaller than the sum of LiDAR weights ($106.1<149.9$). In other words, the learned fusion weights represent overall more reliance on LiDAR than camera, which aligns with the overall better performance of LiDAR-only models over camera-only models. Still, we do observe quite diverse weight values. Certainly not all channels favor LiDAR, and few weights are close to $0.5$, the default for regular average fusion. The general higher influence of LiDAR on the fused results may also explain why camera-only inference of UniBEV performs marginally worse than UniBEV_C, while the LiDAR-only inference of UniBEV even slightly outperforms UniBEV_L.
@@
TABLE II: Comparison of different fusion approaches on nuScenes val set for a fixed decoder dimensions of 256. (best/ second best)
@@ IV-C Ablation Study
We here discuss our ablation results for the different fusion modules and the unified BEV queries.
@@ IV-C1 Comparison of different fusion modules
When evaluating the performance across diverse input modalities, the L+C and L-only performances of UniBEV_CNW improve relative to UniBEV_avg, particularly evident in the L-only performance, but the C-only performance sees a decline. We hypothesize that CNW effectively lets the detector head rely more on LiDAR for the final fusion result, impacting its camera-only performance. Overall, the performance gap between CNW and average fusion appears minor, and if such trade-off is favorable for the target application remains a future research.
IV-C2 Unified Queries vs. Separate Queries
Finally, we compare the performance of UniBEV using shared BEV queries $Q$ across modalities, against a variant that learns separate queries for each modality. Table III shows the model with unified queries has a minor edge over its counterpart with separate queries across all three input combinations, as well as in the summary metric ($52.5\%$ vs. $52.2\%$ summary mAP). Interestingly, the model with separate queries matches the performance with UniBEV_L (refer to Table I, $57.8\%$ mAP) while unified queries model can slightly exceed this LiDAR-only trained counterpart. A possible explanation is that shared queries provide a weak interaction between the BEV encoders during training, which facilitates aligning their feature spaces.
TABLE III: Comparison between separate queries design and unified queries design on nuScenes val set. (best/second best)
@@ V CONCLUSIONS
We have presented UniBEV, a multi-modal 3D object detection model designed with missing sensor modality inputs in mind. The experiments demonstrate UniBEV’s higher robustness to missing inputs compared to SotA BEV-based detection methods, BEVFusion and MetaBEV. UniBEV achieves $52.5\%$ mAP on average over all input combinations, significantly improving over the baselines, with BEVFusion averaging at $43.5\%$ mAP, and MetaBEV averaging at $48.7\%$ mAP. Future work can further investigate what properties multi-modal features should exactly possess for robustness, and if UniBEV’s fused BEV features also benefit other tasks, such as BEV map segmentation.

### v3

@@ IV-B Multi-Modal 3D Object Detection
Table I showcases the inference performance of the fusion-based detector on both multi-modal input, as well as single-modality input using the same trained weights.
Multi-modal robustness. The summary metrics of our UniBEV ($58.7\%$ NDS and $52.5\%$ mAP) significantly surpass the baseline methods, indicating that UniBEV is more robust over varying input modalities. On all input modalities, UniBEV outperforms its multi-modal baselines, achieving $68.5\%$ NDS and $64.2\%$ mAP for LiDAR+camera fusion, especially the difference in camera-only performance is notable. Despite employing a more powerful image backbone [37], BEVFusion lags markedly behind UniBEV when only camera input is available.
Given that the CenterPoint head of BEVFusion and our detection head exhibit comparable detection capabilities (as evidenced by the near identical performance of UniBEV_L and CenterPoint), the performance difference for camera-only between UniBEV and BEVFusion can be attributed to the quality of BEV features and its fusion strategy. Fig. 3 illustrates that compared to BEVFusion, UniBEV’s camera and LiDAR BEV features more clearly discern similar object locations and that these are better spatially aligned. Besides, BEVFusion utilizes LSS [6] as the camera BEV encoder to project image features into the BEV feature space. This enforces an inductive bias on its camera BEV features not present in its LiDAR BEV features, as exhibited by the hexagon-shaped outline.
Figure 3: Example of the BEV feature maps of different modalities for a single sample. High intensity indicates high variance across channels at that location. UniBEV aligns modalities for object detection, resulting in strong responses at the same locations in both the camera and LiDAR.
@@
Qualitative results. To support the comparison between UniBEV and its multi-modal baselines, we show some qualitative detection results in Fig. 4. For instance, we can see that BEVFusion for camera-only suffers from various false negatives, whereas MetaBEV tends to have more false positives.
@@
Fig. 5 (a) visualizes UniBEV’s $N=256$ normalized CNW weights, sorted from most LiDAR weighted to most camera weighted, and also reports their total summed weights. We observe the sum of camera weights is smaller than the sum of LiDAR weights ($106.1<149.9$). In other words, the learned fusion weights represent overall more reliance on LiDAR than on camera, which aligns with the overall better performance of LiDAR-only models over camera-only models. Still, we do observe quite diverse weight values. Certainly, not all channels favor LiDAR, and few weights are close to $0.5$, the default for regular average fusion. The general higher influence of LiDAR on the fused results may also explain why the camera-only inference of UniBEV performs marginally worse than UniBEV_C, while the LiDAR-only inference of UniBEV even slightly outperforms UniBEV_L.
@@
Inference speed. Finally, we measure the inference speed of all multi-modal methods using both input modalities. Using a Nvidia V100 GPU with a batch size of 1, we find that the average inference speed is 0.7 FPS for BEVFusion, 1.4 FPS for MetaBEV, and 1.6 FPS for our UniBEV. Thus, UniBEV achieves the highest speed of all multi-modal methods by a small margin, indicating its improved performance does not come at the cost of efficiency. We do note that BEVFusion uses a more powerful but slower backbone [3, 37]. Also, UniBEV’s inference speed increases when running on multi-modal input, achieving 2.5 FPS for camera-only, and 3.9 FPS for LiDAR-only.
@@
TABLE II: Comparison of different fusion approaches on nuScenes val set for a fixed decoder dimension of 256. The Modality Dropout strategy is applied to all models. (best/ second best)
@@ IV-C Ablation Study
We here discuss our ablation results for the different fusion modules, the effect of probabilities $p_{L}$ and $p_{C}$ during Modality Dropout, and the unified BEV queries.
@@ IV-C1 Comparison of different fusion modules
When evaluating the performance across diverse input modalities, the L+C and L-only performances of UniBEV_CNW improve relative to UniBEV_avg, particularly evident in the L-only performance, but the C-only performance sees a decline. We hypothesize that CNW effectively lets the detector head rely more on LiDAR for the final fusion result, impacting its camera-only performance. Overall, the performance gap between CNW and average fusion appears minor, and if such a trade-off is favorable for the target application remains a future research.
IV-C2 Effect of probabilities $p_{L}$ and $p_{C}$ during Modality Dropout
Next, we investigate the influence of the probabilities of keeping the different modalities during MD. Table III showcases the performance of the model when changing $p_{L}$ and $p_{C}$ while keeping the probability of dropping a modality fixed to $p_{md}=0.5$.
The L+C performance declines significantly in the two extreme cases where $p_{L}=0$ and $p_{L}=1$. However, in the other cases the L+C performance remains mostly stable, even when LiDAR-only and camera-only probabilities are unbalanced. As expected, the performance of LiDAR-only and camera-only consistently improves as the proportion of their respective inputs used during training increases.
As shown in the table, LiDAR-only mAP increases by $12.8$ percentage points as the probability of LiDAR-only training rises from $0\%$ to $75\%$, but already achieves an mAP of $45.5\%$ even without LiDAR-only inputs during training. Remarkably, training with $100\%$ LiDAR-only during MD decreases both LiDAR-only and L+C performance compared to including $25\%$ camera-only training iterations, indicating that the camera-only input also regularizes the network for LiDAR.
The camera-only performance sees a substantial increase in mAP by $33$ percentage points as the proportion of camera-only training iterations increases from $0\%$ to $100\%$. However, unlike LiDAR, the camera-only performance can only reach $3.0\%$ mAP without any training with camera-only inputs, and does not benefit from adding LiDAR-only training iterations.
These observations further support our insights, namely that fusion mostly relies on the more informative sensor, in this case LiDAR, which allows to train the LiDAR features through fusion even with few LiDAR-only training iterations. The same is not true for the camera features, which as the weaker modality strictly relies on MD to make the network achieve good performance. This observation also suggests that especially emphasizing the weaker modality during training could enhance the overall robustness of the model, even for the other modality. Due to the optimal summary and L+C performance, we keep $p_{md}=0.5$ and $p_{L}=p_{C}=0.5$ as the default for all our other experiments. We leave studying the impact of varying the MD probability $p_{md}$ as future work.
TABLE III: Effect of sensor dropping probabilities on nuScenes val set. The MD probability $p_{md}$ is $0.5$. (best/second best/default setup)
mAP
$p_{L}$ $p_{C}$ L+C $\uparrow$ L $\uparrow$ C $\uparrow$ Summary Metric $\uparrow$
0 1 63.2 45.5 36.0 48.2
0.25 0.75 64.0 57.8 35.8 52.5
0.50 0.50 64.2 58.2 35.0 52.5
0.75 0.25 63.8 58.3 33.2 51.8
1 0 60.8 55.9 3.0 39.9
IV-C3 Unified Queries vs. Separate Queries
Finally, we compare the performance of UniBEV using shared BEV queries $Q$ across modalities against a variant that learns separate queries for each modality. Table IV shows the model with unified queries has a minor edge over its counterpart with separate queries across all three input combinations, as well as in the summary metric ($52.5\%$ vs. $52.2\%$ summary mAP). Interestingly, the model with separate queries matches the performance with UniBEV_L (refer to Table I, $57.8\%$ mAP) while the unified queries model can slightly exceed this LiDAR-only trained counterpart. A possible explanation is that shared queries provide a weak interaction between the BEV encoders during training, which facilitates aligning their feature spaces. We conclude that the overall performance difference is only small, though the unified query design has the additional advantage that it requires fewer model parameters. For these reasons, we have adopted it as our default configuration.
TABLE IV: Comparison between separate queries design and unified queries design on nuScenes val set. (best/second best)
@@ V CONCLUSIONS
We have presented UniBEV, a multi-modal 3D object detection model designed with missing sensor modality inputs in mind. The experiments demonstrate UniBEV’s higher robustness to missing inputs compared to SotA BEV-based detection methods, BEVFusion, and MetaBEV. UniBEV achieves $52.5\%$ mAP on average over all input combinations, significantly improving over the baselines, with BEVFusion averaging at $43.5\%$ mAP, and MetaBEV averaging at $48.7\%$ mAP. Our proposed CNW fusion approach demonstrates superior performance compared to the commonly employed concatenation. The analysis of the learned weights reveals the inherent characteristics of a fusion process, notably that the model exhibits a greater reliance on LiDAR features as compared to camera features.
Future research can address dynamically adjusting channel weights, for instance based on the environmental conditions or content of the scene. Another open question remains what properties multi-modal features should exactly possess for robustness. Finally, we will explore if UniBEV’s fused BEV features also benefit other tasks, such as BEV map segmentation. We hope that our work will inspire further research on robust perception for autonomous driving.